OpenStudy (anonymous):

Which equation would best help solve the following problem? The height of a triangle is 6 m more than its base. The area of the triangle is 56 m2. Find the length of the base. prntscr.com/3my8i1

OpenStudy (anonymous):

that m^2 btw

OpenStudy (anonymous):

that is...>.<

OpenStudy (loser66):

can't open the link

OpenStudy (wolf1728):

Try this link: http://prntscr.com/3my8i1

OpenStudy (anonymous):

Thanks @wolf1728

OpenStudy (wolf1728):

Thanks Dragon :-)

OpenStudy (loser66):

why don't you help him @wolf1728

OpenStudy (wolf1728):

I was helping

OpenStudy (loser66):

From the options, I assume that the triangle is a right triangle so if x is the base--> the high is x +6 Area = \(\dfrac{1}{2}height *base\) plug all in to get the answer

OpenStudy (wolf1728):

I'm NOT assuming it is a right triangle.

OpenStudy (anonymous):

I don't know what triangle it is..

OpenStudy (loser66):

ok, I am sorry for my wrong guidance. hihihi...let wait for Wolf

OpenStudy (anonymous):

I have no idea how loser66 got a 99 score

OpenStudy (anonymous):

He is good, just messing around now.

OpenStudy (wolf1728):

I'm trying to draw a graphic Anyway I have a formula ( I think) 56 m² = 1/2 base² + 3*base

OpenStudy (wolf1728):

Thanks for the compliment Dragon "He is good"

OpenStudy (anonymous):

Ok, :) Thanks for helping

OpenStudy (wolf1728):

Okay - I trust you chose the answer from my equation. (I didn't think I was finished) :-)

OpenStudy (loser66):

@jinxhead20 do you know why I got 99 ?? hehehe... I cheated.

OpenStudy (loser66):

go to my detail profile and see, you can see that I pick the very easy one to help, so that everybody got my help gave me medal, and....the result is I got 99 try to get that score and you can see that is no meaning. To me, it's fake. and I know it. hehehe...

OpenStudy (loser66):

@Wolf1728, no matter what the triangle is, the area of it still be 1/2 base *height